One-half of Heather's age two years from now plus one-third of her age three years ago is twenty years. How old is she now?


1
Expert's answer
2021-08-15T18:17:51-0400


Let x be the age of Heather now. Consider and solve the equation:



"\frac{x+2}{2}+\frac{x-3}{3}=20""3(x+2)+2(x-3)=120""3x+6+2x-6=120""5x=120""x=24"

Answer. 24 years.